# Long-Wave Infrared Spintronic Poisson Bolometers with High Sensitivity

### v1

@@ 2 Results
Figure 4 (b) shows the NEDT result measured with dT = 0.3 K. For this applied field and bias voltage labeled on the plot, the device shows a linear response between 24.7 and 25.3 °C. Afterward, the count rate stops increasing due to device saturation. Figure 4 (c) shows a repeated NEDT measurement with dT = 0.1 K and the same applied field and bias voltage as (b), but with a focus on the linear response region to achieve the best possible sensitivity. The result shows an improved NEDT of 35 mK. Figure 4 (d) shows an NEDT measurement with decreased bias voltage and slightly increased applied field. The result shows a decreased count rate and higher NEDT due to both current-induced spin-transfer torque effect and the field dependence of the count rate [bauer2025exploiting]. With the bias voltage and the applied field properly selected, the spintronic Poisson bolometer is able to achieve close to, well below 100 mK NEDT in multiple measurements on different days. Figure 4 (e)-(f) show NEDT measured with negative bias voltage. The results show the tunability of device performance with the bias voltage and applied field. Each of these 6 measurements is done on a different day. The integration time for (b) - (d) is 20 ms (50 Hz frame rate). The integration time for (e) and (f) is 40 ms (25 Hz frame rate), all comparable to the frame rate of state-of-the-art commercial LWIR cameras.
We can normalize the NEDT results with different f-numbers to f/1.0 using:
@@
Table 1 shows a comparison of the normalized NEDT results of the FLIR A325sc LWIR camera and the spintronic Poisson bolometer:
@@
Due to enhanced absorption of thermal radiation, the energy in the sensing layer quickly surpasses the energy barrier, causing the count rate to reach saturation. Beyond this point, further temperature increases no longer produce a significant rise in count rate. Consequently, the spintronic Poisson bolometer demonstrates a limited dynamic range, within which a linear response to the blackbody temperature can be observed. Nevertheless, the sensitivity within this linear regime rivals that of a commercial uncooled LWIR camera.
@@ 3 Conclusion
The Poisson bolometer operates in a probabilistic regime dominated by Poissonian noise, establishing a completely new detection paradigm. In Poisson bolometers, random thermal fluctuations manifest as Poissonian countable events rather than a continuous current or voltage, such as in traditional analog bolometers, allowing the signal and the noise to be separated more effectively, thereby alleviating sensitivity limitations. In this work, a broadband LWIR plasmonic absorber was designed and integrated onto a spintronic Poisson bolometer to enhance LWIR absorption and improve thermal coupling of the incident radiation and the sensing layer. Through this innovation of integrating spintronic materials and plasmonic materials, the device achieved a best NEDT of 35 mK at a 50 Hz frame rate, with multiple measurements close to or below 100 mK, demonstrating room-temperature performance among the most sensitive uncooled LWIR detectors reported to date. These results represent a significant advancement in the development of uncooled infrared detection technologies, approaching the sensitivity traditionally associated with cryogenic systems. The demonstrated performance highlights the potential of this approach for a wide range of high-sensitivity LWIR applications, including remote thermal monitoring, high-speed imaging, cryogenic system diagnostics, and space-based infrared sensing.
@@ Acknowledgements
Please insert your acknowledgments here.

### v3

@@ 2 Results
Figure 4 (b) shows the NEDT result measured with dT = 0.3 K. For this applied field and bias voltage labeled on the plot, the device shows a linear response between 24.7 and 25.3 °C. Afterward, the count rate stops increasing due to device saturation. Figure 4 (c) shows a repeated NEDT measurement with dT = 0.1 K and the same applied field and bias voltage as (b), but with a zoom-in on the linear response region to achieve the best possible sensitivity. The result shows an improved NEDT of 35 mK. Figure 4 (d) shows an NEDT measurement with decreased bias voltage and slightly increased applied field. The result shows a decreased count rate and higher NEDT due to both current-induced spin-transfer torque effect and the field dependence of the count rate [3]. With the bias voltage and the applied field properly selected, the spintronic Poisson bolometer can achieve close to, well below 100 mK NEDT in multiple measurements on different days. Figure 4 (e)-(f) show NEDT measured with negative bias voltage. The results show the tunability of device performance with the bias voltage and applied field. Each of these 6 measurements is done on a different day. The integration time for (b) - (d) is 20 ms (50 Hz frame rate). The integration time for (e) and (f) is 40 ms (25 Hz frame rate), all comparable to the frame rate of state-of-the-art commercial LWIR cameras.
The variation of NEDTs in Figure 4 (b)-(f) is likely caused by two factors. First, both the current-induced spin-transfer torque and the field dependence of the count rate affect the device sensitivity [3]. Second, the response temperature range of the device is currently limited, and measurements with larger temperature steps (dT) generally yield worse sensitivity, as some temperatures fall outside the linear response range. We also notice that the count rate plateaus at higher temperatures, which could result from device saturation and the device’s multi-state behavior. Enhanced absorption of thermal radiation causes the energy in the sensing layer to quickly surpass the energy barrier, leading the count rate to saturate. Beyond this point, further temperature increases no longer produce a significant rise in the count rate. Meanwhile, although the device is dominated by two meta-stable magnetization states, multiple minor states coexist simultaneously. The counting algorithm treats the response as strictly two-state and may overlook increased transitions between the less-frequent states. Consequently, the spintronic Poisson bolometer demonstrates a limited response temperature range, within which a linear response to the blackbody temperature can be observed. Nevertheless, the sensitivity within this linear regime rivals that of a commercial uncooled LWIR camera.
Table 1 shows a comparison of the normalized NEDT results of the VOx microbolometer-based LWIR camera and the spintronic Poisson bolometer. The NEDT results with different f-numbers are normalized to f/1.0 using [43]:
@@
Many high-sensitivity uncooled LWIR detectors are characterized in vacuum chambers [10, 2, 6], which effectively reduces thermal conductance and thereby maximizes the temperature rise induced by incident radiation. We expect that the sensitivity presented in this work could be further enhanced under similar vacuum conditions. The sensitivity achieved in this work establishes spintronic Poisson bolometers as a promising platform for practical applications, such as neuromorphic computing, event-based sensing, and thermal spectral imaging [35, 34, 33]. Combined with emerging photonic and nanoscale detector technologies, these approaches could enable compact, energy-efficient infrared imaging systems suitable for remote sensing, autonomous navigation, industrial inspection, and rapid thermal event detection. More broadly, the methods presented here may inform future designs of statistical sensing systems beyond infrared detection, including other spectral regimes and quantum-limited measurements.
@@ 3 Conclusion
The Poisson bolometer operates in a probabilistic regime dominated by Poissonian noise, establishing a completely new detection paradigm. In Poisson bolometers, random thermal fluctuations manifest as Poissonian countable events rather than as a continuous current or voltage, as seen in traditional analog bolometers, allowing the signal and noise to be separated more effectively, thereby alleviating sensitivity limitations. In this work, a broadband LWIR plasmonic absorber was integrated onto a spintronic Poisson bolometer to enhance LWIR absorption and improve thermal coupling of the incident radiation and the sensing layer. Through this innovation of integrating spintronic materials and plasmonic materials, the device achieved a best NEDT of 35 mK at a 50 Hz frame rate, with multiple measurements close to or below 100 mK, demonstrating room-temperature performance comparable to the most sensitive uncooled LWIR detectors reported to date. These results represent a significant advancement in the development of uncooled infrared detection technologies, approaching the sensitivity traditionally associated with cryogenic systems. The demonstrated performance highlights the potential of this approach for a wide range of high-sensitivity LWIR applications, including remote thermal monitoring, high-speed imaging, cryogenic system diagnostics, and space-based infrared sensing.
@@ Acknowledgements
This work is partially supported by the Elmore Chaired Professorship of Purdue University.
Conflict of Interest
The authors declare no conflict of interest.